OpenStudy (anonymous):

The square root of (7 + 4 root 3)

OpenStudy (saifoo.khan):

\[\Large \sqrt{7+4\sqrt{3}}\]

OpenStudy (anonymous):

\[\sqrt{7+4\sqrt{3}}\] ^that's what it looks like.

OpenStudy (saifoo.khan):

Mine is bigger! :O

OpenStudy (saifoo.khan):

\[2+\sqrt{3} \]

OpenStudy (anonymous):

Thanks, can you tell me how you got that? :)

OpenStudy (anonymous):

The procedure is here:- http://en.wikipedia.org/wiki/Nested_radical#Denesting_nested_radicals